OpenStudy (fanduekisses):

Related Rated: The Volume of a cube is decreasing at a constant rate of 10 cubic centimeters per minute. What is the rate of change, in square centimeters per minute of the surface area of the cube at the instant when the length of each edge of the cube id 5 cm?

OpenStudy (fanduekisses):

So will I first have to find a relationship between the surface area and the volume of a cube?

OpenStudy (anonymous):

\[\prod69_{t}^{h}.dsov rot =21\]

OpenStudy (fanduekisses):

what?

OpenStudy (anonymous):

wait math are u in

OpenStudy (anonymous):

lol srry

OpenStudy (anonymous):

lol cant remember im in cal bc v4

OpenStudy (anonymous):

sorry

OpenStudy (surry99):

you need expressions for the volume of the cube and surface area of the cube to start

OpenStudy (fanduekisses):

\[v=x^{3}\] \[SA= 6x^{2}\] x= edges in a cube

OpenStudy (fanduekisses):

\[\frac{dv}{dt}=x^2*\frac{dx}{dt}\] \[\frac{da}{dt}=12*x*\frac{dx}{dt}\]

OpenStudy (surry99):

yes...so from the first equation you can get dx/dt and plug that in to the second to get da/dt

OpenStudy (fanduekisses):

ok let me try that ^_^

OpenStudy (fanduekisses):

yay I got -8 cm^2 ^_^

OpenStudy (surry99):

-8 cm^2/min

OpenStudy (fanduekisses):

Yeah I forgot about the time :) thanks so much ^_^

OpenStudy (surry99):

you are very welcome!